OpenStudy (anonymous):

Verify the identity. (1 + tan^2u)(1 - sin^2u) = 1

OpenStudy (anonymous):

tan^2u = (sin^2u)/(cos^2u) sin^2u + cos^2u = 1 now what can you do with the above?

OpenStudy (anonymous):

i'm not sure.

OpenStudy (anonymous):

from: sin^2u + cos^2u = 1 you get: cos^2u = 1 - sin^2u how about now?

OpenStudy (anonymous):

i'm sorry i just don't understand this.

OpenStudy (anonymous):

so you know the following identities: tan^2u = (sin^2u)/(cos^2u) cos^2u = 1 - sin^2u substitute the above accordingly to the equation: (1 + tan^2u)(1 - sin^2u) = (1 + (sin^2u)/(cos^2u))(cos^2u) = cos^2u + cos^2u[ (sin^2u)/(cos^2u)] = cos^2u + sin^2u =?

OpenStudy (anonymous):

\[(1+\frac{ \sin^2x }{ \cos^2x })(\cos^2x)=1\]\[\cos^2x+\sin^2x=1\]

OpenStudy (anonymous):

@ChmE How did you skip to that last step? I don't understand it, and I don't think the asker would either, can you please clarify?

OpenStudy (anonymous):

distributed cos^2(x) to each term